# Black-Box $k$-to-$1$-PCA Reductions: Theory and Applications

### v2

@@ 1.2 Our techniques
In other words, as long as we can appropriately control the quantity $\lVert\mathbf{U}_{1}^{\top}\mathbf{L}\bm{\Lambda}\mathbf{L}^{\top}\widetilde{\mathbf{L}}\rVert_{\textup{op}}$, we can bound the parameter degradation (from $(\delta,\gamma)$ to $(\Delta,\Gamma)$) of composing two cPCAs.
@@ 1.3 Related work
We are aware of relatively few papers considering deflation-based approaches to $k$-PCA without strong gap assumptions on the target matrix (see e.g. [LZ15] for a gap-based analysis). Perhaps the most direct comparison is [AZL16], who analyze deflation algorithms in the white-box setting, where the target matrix is explicit. The result of [AZL16] are incomparable to ours, as it is reliant on the ability to query matrix-vector products to solve linear systems to high precision using first-order methods, where precision levels scale roughly as $\epsilon^{\Omega(\log(d))}$, for an approximation parameter $\epsilon$. We view the [AZL16] result as complementary to our Theorems 1 and 2, which analyze deflation methods under minimal assumptions on the $1$-PCA oracle beyond its approximation guarantees.
We also draw inspiration from the extensive body of work focusing on computationally efficient and statistically optimal algorithms for PCA, e.g. [MM15, ${}^{+}$16, Sha16b] which give runtime bounds in the white-box setting, and e.g. [HP14, SRO15, BDWY16, Sha16a, ${}^{+}$16, LWLZ18] which focus on offline and online variants of statistical PCA. In particular, the gap-free notion of approximation in our cPCA definition is patterned off analogous definitions in this latter line.
Additionally, we mention the contemporaneous work of [MMM24], which gives analyses of black-box $k$-to-$1$ reductions for block Krylov methods targeting low-rank approximation, which is very close in spirit to our work. However, [MMM24] again requires explicit access to matrix-vector products for their reductions, which makes their results not directly applicable in our setting.
@@ 2 Preliminaries
For a subspace $V\subseteq\mathbb{R}^{d}$, we let $\mathbf{P}_{V}\in\mathbb{R}^{d\times d}$ be the orthogonal projection matrix onto $V$. We often identify a subspace $V\subseteq\mathbb{R}^{d}$ with a generating matrix $\mathbf{V}\in\mathbb{R}^{d\times r}$, where $r:=\dim(V)$, such that $\mathbf{V}$ has orthonormal columns and $\operatorname{span}(\mathbf{V})=V$. In this case, note that $\mathbf{P}_{V}=\mathbf{V}\mathbf{V}^{\top}$ for any valid $\mathbf{V}$. We write $U\perp V$ for subspaces $U,V\subset\mathbb{R}^{d}$ if every element of $U$ is orthogonal to every element of $V$. Similarly, for two orthonormal matrices $\mathbf{U},\mathbf{V}$ with the same number of rows, we write $\mathbf{U}\perp\mathbf{V}$ if $\operatorname{span}(\mathbf{U})\perp\operatorname{span}(\mathbf{V})$. For orthonormal $\mathbf{U}\in\mathbb{R}^{d\times r}$, we write $\mathbf{U}_{\perp}$ to mean any orthonormal $d\times(d-r)$ matrix so that columns of $\mathbf{U}$ and $\mathbf{U}_{\perp}$ form a basis of $\mathbb{R}^{d}$, and define $U_{\perp}$ to be the orthogonal complement of a subspace $U\subseteq\mathbb{R}^{d}$. For $\mathbf{M}\in\mathbb{S}^{d\times d}$ and $\lambda>0$, we denote the subspace spanned by eigenvectors of $\mathbf{M}$ corresponding to eigenvalues at least $\lambda$ by $V^{\geq\lambda}(\mathbf{M})$, and the corresponding orthonormal generating matrix by $\mathbf{V}^{\geq\lambda}(\mathbf{M})$. We similarly define $V^{>\lambda}(\mathbf{M})$, $V^{\leq\lambda}(\mathbf{M})$, etc. Finally, we provide a few PCA-related definitions used throughout the paper.
Definition 4 (Exact PCA).
For $\mathbf{M}\in\mathbb{S}_{\succeq\mathbf{0}}^{d\times d}$, the exact $k$-principal component analysis ($k$-PCA) problem asks to return orthonormal $\mathbf{V}^{\geq\lambda_{k}(\mathbf{M})}(\mathbf{M})$ if it is $k$-dimensional, or more generally (i.e. in the case when $\lambda_{k+1}(\mathbf{M})=\lambda_{k}(\mathbf{M})$) any orthonormal $\mathbf{V}\in\mathbb{R}^{d\times k}$ such that
$V^{\geq\lambda_{k}(\mathbf{M})}(\mathbf{M})\supseteq\operatorname{span}(\mathbf{V})\supset V^{>\lambda_{k}(\mathbf{M})}(\mathbf{M}).$
We call such $\mathbf{V}$ an exact $k$-PCA of $\mathbf{M}$.
Definition 5 (ePCA oracle).
Definition 6 (cPCA oracle).
We say $\mathcal{O}_{\textup{cPCA}}$ is a $(\delta,\gamma)$-approximate $1$-cPCA oracle (or, $(\delta,\gamma)$-$1$-cPCA oracle) if, on inputs $\mathbf{M}\in\mathbb{S}_{\succeq\mathbf{0}}^{d\times d}$, and $\mathbf{P}\in\mathbb{R}^{d\times d}$, where $\mathbf{P}$ is required to be an orthogonal projection matrix, $\mathcal{O}_{\textup{cPCA}}$ returns $u\in\mathbb{R}^{d}$, a $(\delta,\gamma)$-$1$-cPCA of $\mathbf{P}\mathbf{M}\mathbf{P}$, satisfying $u\in\operatorname{span}(\mathbf{P})$.
@@
Finally, we make the following basic observations relating Definitions 1 and 2 which allow us to convert an ePCA guarantee to a cPCA one and vice-versa.
Lemma 1.
@@
It is important to note that naïvely applying such conversions with our lossless ePCA result in Theorem 1 do not provide us with optimal cPCA overheads, as in Theorem 2, even when $k=2$ and $\kappa_{k}(\mathbf{M})=1$. For example, consider the strategy which defines, for some $\Delta,\Gamma\in(0,1)$, $\epsilon:=\Delta\Gamma$, and then uses Theorem 1 to obtain an $\epsilon$-$2$-ePCA by calling an $\epsilon$-approximate $1$-ePCA oracle twice. By Lemma 1, the result is indeed a $(\Delta,\Gamma)$-$2$-cPCA of $\mathbf{M}$. However, for a $(\delta,\gamma)$-approximate $1$-cPCA oracle to be an $\epsilon$-approximate ePCA oracle (via Lemma 2), this requires taking $\delta,\gamma\leq\epsilon=\Delta\Gamma$. Finally, standard analyses of the number of samples required to implement such a $(\delta,\gamma)$-$1$-cPCA oracle (see [${}^{+}$16]), in the standard, i.i.d. setting, scale as $\frac{1}{\delta\gamma^{2}}=\frac{1}{\Delta^{3}\Gamma^{3}}$, which is polynomially worse than the sample complexity of $\frac{1}{\Delta\Gamma^{2}}$ required to directly obtain a $(\Delta,\Gamma)$-cPCA in this setting. Our Theorem 2 does not suffer from such lossiness, up to an overhead only depending on $k$.
@@ 4.2 Invalid regimes for black-box cPCA
In this section, we give impossibility results which establish several parameter ranges on $(\Delta,\Gamma=g(\Delta,\kappa_{k}(\mathbf{M})))$, characterized by a function $g:[0,1]\times(1,\infty)\to[0,1]$, which fall in the invalid black-box cPCA regime (Definition 3). We first show that sublinear functions induce invalid regimes.

### v3

@@ 1.2 Our techniques
In other words, as long as we can appropriately control the quantity $\lVert\mathbf{U}_{1}^{\top}\mathbf{L}\bm{\Lambda}\mathbf{L}^{\top}\widetilde{\mathbf{L}}\rVert_{\textup{op}}$, we can bound the parameter degradation (from $(\delta,\gamma)$ to $(\Delta,\Gamma)$) of composing two cPCAs. For instance, if this quantity is $O(\gamma\lambda_{k}(\mathbf{M}))$, then cPCA composition blows up both the $\delta$ and $\gamma$ parameters by constants.
@@ 1.3 Related work
We are aware of few papers considering deflation-based approaches to $k$-PCA without strong gap assumptions on the target matrix (see e.g. [LZ15] for a gap-based analysis). The most direct comparison is [AZL16], who analyze Algorithm 1 when $\mathcal{O}_{1\textup{PCA}}$ is a $1$-cPCA oracle, giving guarantees summarized in (1). Our Theorem 1 shows the lossiness in the [AZL16] reduction is unnecessary if our approximations are instead measured through Definition 2 (ePCA), and yields a competitive guarantee (3) for cPCA. Our Theorem 2 further removes all overhead in this cPCA reduction for $k=\Theta(1)$, in the valid parameter regimes not ruled out by our lower bound, Proposition 1.
We also draw inspiration from the extensive body of work focusing on computationally efficient and statistically optimal algorithms for PCA, e.g. [MM15, GHJ+16, Sha16] which give runtime bounds in the white-box setting, and e.g. [HP14, SRO15, BDWY16, Sha16, JJK+16, LWLZ18] which focus on offline and online variants of statistical PCA. In particular, the gap-free notion of approximation in our cPCA definition is patterned off analogous definitions in this latter line. Further, several other notions of PCA approximation have been proposed in the white-box setting, e.g. Eqs. (1)-(3) of [MM15]; however, all such results we are aware of use matrix-vector products. Our motivation is analyzing Algorithm 1 as a reduction, so we focused on error metrics where black-box $1$-PCA guarantees already exist, though it is interesting to prove similar reductions for other metrics.
Additionally, we mention the contemporaneous work of [MMM24], which gives analyses of variants of $k$-to-$1$ reductions for block Krylov methods targeting low-rank approximation, which is very close in spirit to our work. However, [MMM24] again requires access to matrix-vector products for their reductions, which makes their results not directly applicable in our black-box setting.
@@ 2 Preliminaries
For a subspace $V\subseteq\mathbb{R}^{d}$, we let $\mathbf{P}_{V}\in\mathbb{R}^{d\times d}$ be the orthogonal projection matrix onto $V$. We often identify a subspace $V\subseteq\mathbb{R}^{d}$ with a generating matrix $\mathbf{V}\in\mathbb{R}^{d\times r}$, where $r:=\dim(V)$, such that $\mathbf{V}$ has orthonormal columns and $\operatorname{span}(\mathbf{V})=V$. In this case, note that $\mathbf{P}_{V}=\mathbf{V}\mathbf{V}^{\top}$ for any valid $\mathbf{V}$. We write $U\perp V$ for subspaces $U,V\subset\mathbb{R}^{d}$ if every element of $U$ is orthogonal to every element of $V$. Similarly, for two orthonormal matrices $\mathbf{U},\mathbf{V}$ with the same number of rows, we write $\mathbf{U}\perp\mathbf{V}$ if $\operatorname{span}(\mathbf{U})\perp\operatorname{span}(\mathbf{V})$. For orthonormal $\mathbf{U}\in\mathbb{R}^{d\times r}$, we write $\mathbf{U}_{\perp}$ to mean any orthonormal $d\times(d-r)$ matrix so that columns of $\mathbf{U}$ and $\mathbf{U}_{\perp}$ form a basis of $\mathbb{R}^{d}$, and define $U_{\perp}$ to be the orthogonal complement of a subspace $U\subseteq\mathbb{R}^{d}$. For $\mathbf{M}\in\mathbb{S}^{d\times d}$ and $\lambda>0$, we denote the subspace spanned by eigenvectors of $\mathbf{M}$ corresponding to eigenvalues at least $\lambda$ by $V^{\geq\lambda}(\mathbf{M})$, and the corresponding orthonormal generating matrix by $\mathbf{V}^{\geq\lambda}(\mathbf{M})$. We similarly define $V^{>\lambda}(\mathbf{M})$, $V^{\leq\lambda}(\mathbf{M})$, etc.
@@
Finally, we prove the following basic observations relating Definitions 2 and 4, stated earlier, which allow us to convert an ePCA guarantee to a cPCA one and vice-versa.
See 1
@@
It is important to note that naïvely applying such conversions with our lossless ePCA result in Theorem 1 do not provide us with optimal cPCA overheads, as in Theorem 2, even when $k=2$ and $\kappa_{k}(\mathbf{M})=1$. Our Theorem 2 does not suffer from such lossiness, up to an overhead only depending on $k$, except in a parameter regime ruled out by Proposition 1.
@@ 4.2 Invalid regimes for black-box cPCA
Before we prove our main result on cPCA (Theorem 2) in the following Section 4.3, in this section, we point out that not all values of $\left(\delta,\gamma\right)$ pairs are feasible inputs for Algorithm 1 to serve as a lossless black-box reduction even for constant $k$, unlike in the case of ePCA. Formally, we characterize two types of parameter regimes for black-box cPCA using the following definition.
Definition 7 (Black-box cPCA regimes).
$\delta:=\frac{\Delta}{f(k)},\;\gamma:=\frac{\Gamma}{f(k)},$ (15)
Our choice to parameterize the regimes $g$ in Definition 7 by $\kappa_{k}(\mathbf{M})$ is motivated by a result of [Lia23] (see Eq. (1.3) in that paper), which proves that some sample complexity overhead in $\kappa_{k}(\mathbf{M})$ is necessary in general. On the other hand, the requirement that $\Gamma\gtrsim\sqrt{\Delta}$ appears to be an artifact of the specific deflation strategy in Algorithm 1. We now state our main impossibility result.
Proposition 1 (Invalid black-box cPCA regimes).
Moreover, Theorem 2 shows that the complement of the regimes ruled out by Proposition 1 are all valid regimes, where we suffer parameter degradation only in $k$. Hence, Proposition 1 and Theorem 2 completely characterize valid and invalid black-box cPCA regimes, under Definition 7.
We establish Proposition 1 by giving $3$-dimensional examples where Algorithm 1 fails to return a $(\Delta,\Gamma)$-$2$-cPCA if $\Gamma\lesssim\kappa_{2}(\mathbf{M})\sqrt{\Delta}$ asymptotically, assuming $\mathcal{O}_{1\textup{PCA}}$ is a $(\delta,\gamma)$-$1$-cPCA oracle, for any $\delta,\gamma$ smaller than $\Delta,\Gamma$ by a constant. First, we show that sublinear functions induce invalid regimes.